# Text-controlled Motion Mamba: Text-Instructed Temporal Grounding of Human Motion

### v1

@@ Abstract.
Human Motion Analysis, Temporal Grounding, Mamba
journalyear: 2024copyright: acmcopyrightconference: Proceedings of the 32st ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, Australiabooktitle: Proceedings of the 32st ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, Australiaprice: 15.00doi: XXXXXXX.XXXXXXXisbn: 978-x-xxxx-xxxx-x/YY/MMsubmissionid: 1147ccs: Computing methodologies Computer vision
Figure 1. Illustration of the Text-based Human Motion Grounding (TMHG) task and samples of the proposed BABEL-Grounding dataset. Best viewed in color.
1. Introduction
@@
Human motion understanding is a crucial task with a wide range of applications. Recent years have witnessed the flourishing of large-scale motion capture databases (Mahmood et al., 2019; Shahroudy et al., 2016; Mandery et al., 2015; Ionescu et al., 2013), which greatly facilitate the end-to-end training of various motion-related tasks. Based on these databases, recent studies (Plappert et al., 2016; Guo et al., 2022a; Liang et al., 2023; Han et al., 2023b; Tang et al., 2023; Li et al., 2023) have augmented motion datasets with textual annotations. These annotated motion-text pairs enable a range of text-motion tasks that require a joint understanding of human motion and language, such as text-based motion generation (Tevet et al., 2022; Petrovich et al., 2022; Zhang et al., 2023c, 2024a), motion captioning (Guo et al., 2022b; Jiang et al., 2024), text-based motion editing (Goel et al., 2023; Kim et al., 2023; Zhang et al., 2023b), and motion question answering (Endo et al., 2023). However, in real-world scenarios, semantic actions often occur sparsely within lengthy motion sequences. Thus, a textual description of particular human actions often corresponds to specific temporal segments of the sequence rather than the entire sequence. Precisely localizing the temporal segments corresponding to the text query presents a substantial challenge.
@@
In this study, we introduce the task of text-based human motion grounding (THMG) for the first time, which aims to identify the start and end timestamps of all segments corresponding to a given textual description from an untrimmed motion sequence. Unlike existing motion temporal action localization tasks (Yu and Fujiwara, 2023), the queries in THMG consist of arbitrary natural languages rather than a predefined set of action labels. As shown in Figure 1, the THMG task is highly challenging as it requires simultaneous consideration of several critical factors: (1) Achieving precise grounding of the time interval corresponding to the query within a long sequence demands the model’s ability to grasp global temporal context effectively. (2) The model should jointly tackle the information from motion and language, ensuring a thorough fusion and interaction between the two modalities. (3) As human pose representation inherently possesses a graph structure, the model needs to capture the latent spatial topological information.
@@
Effectively capturing global temporal information is crucial for THMG task. Existing frameworks for text-motion analysis mostly adopt recurrent neural networks (Guo et al., 2020; Wang et al., 2022) or temporal convolutions (Zhang et al., 2023c; Zhou et al., 2023; Jiang et al., 2024; Guo et al., 2022b, a; Zhong et al., 2023) as the main workhorse. However, their ability to capture long-term dependency is quite limited. Recently, there has been a surge of interest in transformer-based models (Petrovich et al., 2022; Tevet et al., 2022; Zhai et al., 2023) for modeling temporal dependencies. In these approaches, a global temporal self-attention mechanism is employed across all frames in the sequence. Nonetheless, these methods encounter challenges when dealing with untrimmed sequences that are very long, as computing a global temporal self-attention is exceedingly computationally expensive in such scenarios.
@@
In this work, we aim to attack all the aforementioned challenges through a unified model that seamlessly incorporates temporal global context, language query control, and spatial graph topology. Our primary source of inspiration stems from the recently proposed state space model called Mamba (Gu and Dao, 2023), an efficient model for handling long-term dependencies within lengthy sequences, while maintaining linear computational cost. Mamba has demonstrated its power across diverse domains, including language modeling and visual understanding. However, its potential application in human motion tasks remains largely unexplored.
@@
The core of Mamba is an input-dependent selection mechanism, enabling the model to selectively propagate or forget information over time depending on the current input. This innovative design greatly enhances traditional State Space Model (SSM) methods, as Mamba is capable of grasping the global context of long sequences while filtering out irrelevant information. However, in the THMG task, the model must select information based textual queries as well. To address this challenge, a text-controlled selection mechanism is introduced, wherein the key idea is to condition the state transition matrix on both motion and text queries. Unlike existing multimodal Mamba methods (Qiao et al., 2024; Zhao et al., 2024) that merely concatenate the textual and visual features and feed it into Mamba blocks, our approach is the first work in enabling texts to dictate the selective propagation of input information. This ensures that the model dynamically adjusts its focus based on the interplay between motion and text inputs. Furthermore, as human motion sequences inherently possess a graph topology, the original Mamba model is not suitable as it is designed to operate on univariate time series. To address this, we enhance Mamba by integrating relational information via graph neural networks into its state representation to facilitate topology awareness. The resulting framework, termed Text-Controlled Motion Mamba (TM-Mamba), can selectively extract relevant global context information based on textual queries in the motion sequence.
@@
For evaluation, existing datasets (Punnakkal et al., 2021; Lin et al., 2023b) with frame-level temporal annotations are not directly applicable to the THMG task. While BABEL (Punnakkal et al., 2021) provides temporal boundaries for all actions that occur in the sequence, it lacks complete detailed textual descriptions, offering only simple categorical phrases. Motion-X (Lin et al., 2023b) offers detailed pose descriptions for each frame, but its texts focus on the movement of human body parts at each timestep, making it incapable of establishing a mapping from semantic action descriptions to temporal segments. To address this gap, based on BABEL, a new dataset called BABEL-Grounding is introduced to serve as a benchmark for evaluating THMG task. BABEL-Grounding is the first text-motion dataset that provides detailed textual descriptions of human actions along with their corresponding temporal segments in untrimmed motion sequences. Like real-world scenarios, each query may correspond to multiple temporal segments. Extensive evaluations of TM-Mamba are conducted on the newly introduced dataset, demonstrating its effectiveness in THMG task. Our primary contributions can be summarized as follows:
@@
2. Related work
2.1. Datasets for Text-Motion Learning
@@
This section presents an overview of existing human motion datasets annotated with texts. These text-motion datasets are primarily developed for text-driven motion generation task, hence they typically include textual descriptions at the sequence level for each motion sequence. For example, the KIT Motion Language dataset (Plappert et al., 2016) is the first that provides human motion alongside corresponding sequence-level textual descriptions. Following this direction, several subsequent works have endeavored to construct larger-scale datasets of similar kind, such as HumanML3D (Guo et al., 2022a), InterHuman (Liang et al., 2023), HumanLong3D (Han et al., 2023b), FLAG3D (Tang et al., 2023), STDM (Li et al., 2023). Some datasets go beyond mere textual annotations and incorporate additional contextual information. For instance, HUMANISE (Wang et al., 2022) integrates 3D scene information to facilitate motion generation within 3D environments, while HOI-Diff (Peng et al., 2023) incorporates object geometry information to support human-object interaction during generation. Furthermore, there are datasets tailored for diverse text-motion tasks beyond generation alone. Examples include PoseScript (Delmas et al., 2022), which focuses on static pose generation and pose captioning, and PoseFix (Delmas et al., 2023), which targets motion editing tasks.
@@
However, all of the aforementioned datasets only contain sequence-level annotations, rendering them unsuitable for temporal tasks. To address this limitation, various efforts have been made to develop motion datasets with specific temporal information. For instance, BABEL (Punnakkal et al., 2021) introduces a motion dataset with frame-wise annotations, providing temporal spans for each action label, thereby enabling tasks such as action localization. Constructed upon BABEL, BABEL-QA (Endo et al., 2023) extends the dataset by incorporating question-answer pairs, aiming to facilitate motion-based question answering. Another example is HuMMan-MoGen (Zhang et al., 2023b), which is built upon the HuMMan (Cai et al., 2022) dataset, where each sequence is divided into predefined action phases, along with phase-level detailed annotations describing the movement of each body part. Recently, Motion-X (Lin et al., 2023b) offers part-level textual annotations for human pose at each frame. However, the annotations of Motion-X are based on individual poses and lack a mapping from textual descriptions of semantic actions to temporal boundaries.
2.2. Text-Motion Multi-modal Learning
@@
Recently, there has been a growing interest in text-motion multi-modal learning. Current research mainly focuses on text-to-motion task (also known as text-driven motion generation), where human motion sequences are generated based on natural language (Ahuja and Morency, 2019; Ghosh et al., 2021; Yuan et al., 2023; Ma et al., 2022; Jin et al., 2023; Zhang et al., 2023b; Azadi et al., 2023; Wang et al., 2023; Kong et al., 2023; Karunratanakul et al., 2023; Qian et al., 2023; Li et al., 2023; Zhong et al., 2023; Tevet et al., 2022; Petrovich et al., 2022; Chen et al., 2023; Zhou and Wang, 2023; Zhang et al., 2023d; Lin et al., 2023a; Yang et al., 2023; Zhai et al., 2023; Han et al., 2023b, a; Hoang et al., 2024; Xie et al., 2023b; Zhang et al., 2023a; Shafir et al., 2023; Wei et al., 2023; Xie et al., 2023a; Wang et al., 2022; Guo et al., 2022a; Athanasiou et al., 2022). The key challenge is to learn a joint embedding space for motion and text. For instance, MotionCLIP (Tevet et al., 2022) aligns human motion embeddings with CLIP space through cosine similarity loss to inherit the semantic structure of CLIP latent space. A traditional transformer-based auto-encoder is employed to generate motion sequences. TEMOS (Petrovich et al., 2022) addresses stochastic motion generation using a VAE structure, aligning the embedding space of motion and text via KL-divergence between their latent distributions. Other works focus on employing various modern generation methods for improved conditional generation. For instance, T2M-GPT (Zhang et al., 2023c) utilizes Vector Quantised Variational AutoEncoder (VQ-VAE) to encode motion sequences into discrete tokens, enabling GPT-like autoregressive generation and training through next-token prediction, with text embedding serving as prior. MotionDiffuse (Zhang et al., 2024a) incorporates Denoising Diffusion Probabilistic Models (DDPM) into the task of text-motion generation, greatly enhancing the diversity and fidelity of generated sequences.
@@
In recent years, some works (Dabral et al., 2023; Guo et al., 2022b; Jiang et al., 2024; Zhou et al., 2023; Kim et al., 2023) aim for more general text-motion models capable of seamlessly handling various text-motion tasks simultaneously. For example, TM2T (Guo et al., 2022b) establishes a bi-modal mutual mapping between texts and tokenized human motion using autoregressive neural machine translators (NMT), effectively handling both text-to-motion and motion-to-text tasks. MotionGPT (Jiang et al., 2024) utilizes VQ-VAE to create a motion tokenizer and vocabulary, and then perform pretraining on both motion and text in a unified manner, by treating human motion as a foreign language. Additionally, some works explore various text-motion tasks beyond generation, including motion question answering (Endo et al., 2023), text-based motion editing (Goel et al., 2023; Kim et al., 2023; Zhang et al., 2023b), text-motion retrieval (Petrovich et al., 2023; Messina et al., 2023), etc.
2.3. State Space Model
@@
State space models (SSMs) are a series of sequential models renowned for their computation and memory efficiency and the ability to model long-term dependencies. The pioneering work, S4 (Gu et al., 2021), first proposed applying HiPPO (Gu et al., 2020) initialization to enable SSMs to maintain long-range memory. Subsequent studies (Smith et al., 2022; Fu et al., 2022; Mehta et al., 2022; Gu et al., 2022a; Gupta et al., 2022; Hasani et al., 2022; Gu et al., 2022b) have followed this direction, further improving the space structure and network architecture of S4.
@@
Recently, Mamba (Gu and Dao, 2023) introduces an input-dependent selection mechanism into SSMs, demonstrating linear time efficiency in long-sequence modeling and achieving outstanding performance across various sequential tasks. The model has been adapted to diverse tasks, including image restoration (Guo et al., 2024), image segmentation (Wang and Ma, 2024; Ruan and Xiang, 2024; Ma et al., 2024; Xing et al., 2024; Liu et al., 2024b), point cloud (Liang et al., 2024), video understanding (Yang et al., 2024; Li et al., 2024), pan-sharpening (He et al., 2024), graph analysis (Behrouz and Hashemi, 2024; Wang et al., 2024), multimodal learning (Qiao et al., 2024; Zhao et al., 2024). There’re also some efforts (Zhu et al., 2024; Liu et al., 2024a) aiming to establish a universal visual backbone based on Mamba by sequentializing images using patch-based methods akin to ViT. In a more recent study,  (Zhang et al., 2024b) introduces a motion generation model that incorporates the Mamba block within a denoising U-Net architecture to effectively enhance motion consistency across frames. The textual information is integrated during the diffusion process via a conditional denoiser, while the Mamba block is solely utilized for spatial-temporal feature extraction, without a multimodal design.
@@
Figure 2. Dataset statistics of BABEL-Grounding. ‘Frame Number’ refers to the length of motion sequences. ‘Text Query Length’ denotes the length of textual annotations in the data. ‘Grounded Length Ratio’ indicates the ratio of the length of temporal segments corresponding to each text query to the total length of the sequence. ‘Segment Counts per Query’ refers to the number of temporal segments corresponding to each text query.

### v2

@@ Abstract
I Introduction
@@
Human motion understanding is a crucial task with a wide range of applications. Recent years have witnessed the flourishing of large-scale motion capture databases [1, 2, 3, 4], which greatly facilitate the end-to-end training of various motion-related tasks, such as motion prediction [5, 6, 7, 8], action recognition [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], motion segmentation [21, 22], etc. Based on these databases, recent studies [23, 24, 25, 26, 27, 28] have augmented motion datasets with textual annotations. These annotated motion-text pairs enable a range of text-motion tasks that require a joint understanding of human motion and language, such as text-based motion generation [29, 30, 31, 32], motion captioning [33, 34], text-based motion editing [35, 36, 37], and motion question answering [38]. However, in real-world scenarios, semantic actions often occur sparsely within lengthy motion sequences. Thus, a textual description of particular human actions often corresponds to specific temporal segments of the sequence rather than the entire sequence. Precisely localizing the temporal segments with text query presents a substantial challenge, which has broad applications across domains like surveillance, sports analytics, healthcare and autonomous driving.
In this study, we introduce the task of text-based human motion grounding (THMG) for the first time, which aims to identify the start and end timestamps of all segments corresponding to a given textual description from an untrimmed motion sequence.
Figure 1: Illustration of the Text-based Human Motion Grounding (TMHG) task and samples of the proposed BABEL-Grounding dataset. Best viewed in color.
@@
A closely related task is video-based temporal grounding [39, 40, 41, 42, 43] or video-based action localization [44, 45, 46, 47, 48, 49, 50], which seek to identify the temporal segments within an RGB video that correspond to a given textual description or action class. Yet, in application like gaming, AR, robotics, the RGB modality is unavailable, and data is inherently represented as 3D motion. Also, RGB and 3D motion capture different aspects of human activity: RGB encodes visual appearance, while 3D motion represents kinematics (joints, velocities, rotations), which offers unique benefits like robustness to background, privacy preservation, viewpoint invariance, and lightweight due to reduced dimensionality. Unlike existing motion temporal action localization tasks [51], the queries in THMG consist of arbitrary natural languages rather than a predefined set of action labels. As shown in Figure 1, the THMG task is highly challenging as it requires simultaneous consideration of several critical factors: (1) Achieving precise grounding of the time interval corresponding to the query within a long sequence demands the model’s ability to grasp global temporal context effectively. (2) The model should jointly tackle the information from motion and language, ensuring a thorough fusion and interaction between the two modalities. (3) As human pose representation inherently possesses a graph structure, the model needs to capture the latent spatial topological information.
@@
Effectively capturing global temporal information is crucial for THMG task. Existing frameworks for text-motion analysis mostly adopt recurrent neural networks [52, 53] or temporal convolutions [31, 54, 34, 33, 24, 55] as the main workhorse. However, their ability to capture long-term dependency is quite limited. Recently, there has been a surge of interest in Transformer-based models [30, 29, 56] for modeling temporal dependencies. In these approaches, a global temporal self-attention mechanism is employed across all frames in the sequence. Nonetheless, these methods encounter challenges when dealing with untrimmed sequences that are very long, as computing a global temporal self-attention is exceedingly computationally expensive in such scenarios.
@@
In this work, we aim to attack all the aforementioned challenges through a unified model that seamlessly incorporates temporal global context, language query control, and spatial graph topology. Our primary source of inspiration stems from the recently proposed state space model called Mamba [57], an efficient model for handling long-term dependencies within lengthy sequences, while maintaining linear computational cost. Mamba has demonstrated its power across diverse domains, including language modeling and visual understanding. However, its potential application in human motion tasks remains largely unexplored.
@@
The core of Mamba is an input-dependent selection mechanism, enabling the model to selectively propagate or forget information over time depending on the current input. This innovative design greatly enhances traditional State Space Model (SSM) methods, as Mamba is capable of grasping the global context of long sequences while filtering out irrelevant information. However, in the THMG task, the model must select information based textual queries as well. To address this challenge, a text-controlled selection mechanism is introduced, wherein the key idea is to condition the state transition matrix on both motion and text queries. Unlike existing multimodal Mamba methods [58, 59] that merely concatenate the textual and visual features and feed it into Mamba blocks, our approach is the first work in enabling texts to dictate the selective propagation of input information. This ensures that the model dynamically adjusts its focus based on the interplay between motion and text inputs. Furthermore, as human motion sequences inherently possess a graph topology, the original Mamba model is not suitable as it is designed to operate on univariate time series. To address this, we enhance Mamba by integrating relational information via graph neural networks into its state representation to facilitate topology awareness. The resulting framework, termed Text-Controlled Motion Mamba (TM-Mamba), can selectively extract relevant global context information based on textual queries in the motion sequence.
@@
For evaluation, existing datasets [60, 61] with frame-level temporal annotations are not directly applicable to the THMG task. While BABEL [60] provides temporal boundaries for all actions that occur in the sequence, it lacks complete detailed textual descriptions, offering only simple categorical phrases. Motion-X [61] offers detailed pose descriptions for each frame, but its texts focus on the movement of human body parts at each timestep, making it incapable of establishing a mapping from semantic action descriptions to temporal segments. To address this gap, based on BABEL, a new dataset called BABEL-Grounding is introduced to serve as a benchmark for evaluating THMG task. BABEL-Grounding is the first text-motion dataset that provides detailed textual descriptions of human actions along with their corresponding temporal segments in untrimmed motion sequences. Like real-world scenarios, each query may correspond to multiple temporal segments. Extensive evaluations of TM-Mamba are conducted on the newly introduced dataset, demonstrating its effectiveness in THMG task. Our primary contributions can be summarized as follows:
@@
II Related work
II-A Datasets for Text-Motion Learning
@@
This section presents an overview of existing human motion datasets annotated with texts. These text-motion datasets are primarily developed for text-driven motion generation task, hence they typically include textual descriptions at the sequence level for each motion sequence. For example, the KIT Motion Language dataset [23] is the first that provides human motion alongside corresponding sequence-level textual descriptions. Following this direction, several subsequent works have endeavored to construct larger-scale datasets of similar kind, such as HumanML3D [24], InterHuman [25], HumanLong3D [26], FLAG3D [27], STDM [28]. Some datasets go beyond mere textual annotations and incorporate additional contextual information. For instance, HUMANISE [53] integrates 3D scene information to facilitate motion generation within 3D environments, while HOI-Diff [62] incorporates object geometry information to support human-object interaction during generation. Furthermore, there are datasets tailored for diverse text-motion tasks beyond generation alone. Examples include PoseScript [63], which focuses on static pose generation and pose captioning, and PoseFix [64], which targets motion editing tasks.
@@
However, all of the aforementioned datasets only contain sequence-level annotations, rendering them unsuitable for temporal tasks. To address this limitation, various efforts have been made to develop motion datasets with specific temporal information. For instance, BABEL [60] introduces a motion dataset with frame-wise annotations, providing temporal spans for each action label, thereby enabling tasks such as action localization. Constructed upon BABEL, BABEL-QA [38] extends the dataset by incorporating question-answer pairs, aiming to facilitate motion-based question answering. Another example is HuMMan-MoGen [37], which is built upon the HuMMan [65] dataset, where each sequence is divided into predefined action phases, along with phase-level detailed annotations describing the movement of each body part. Recently, Motion-X [61] offers part-level textual annotations for human pose at each frame. However, the annotations of Motion-X are based on individual poses and lack a mapping from textual descriptions of semantic actions to temporal boundaries.
II-B Text-Motion Multi-modal Learning
@@
Recently, there has been a growing interest in text-motion multi-modal learning. Current research mainly focuses on text-to-motion task (also known as text-driven motion generation), where human motion sequences are generated based on natural language [66, 67, 68, 69, 70, 37, 71, 72, 73, 74, 75, 28, 55, 29, 30, 76, 77, 31, 78, 79, 56, 26, 80, 81, 82, 83, 84, 85, 86, 53, 24, 87]. The key challenge is to learn a joint embedding space for motion and text. For instance, MotionCLIP [29] aligns human motion embeddings with CLIP space through cosine similarity loss to inherit the semantic structure of CLIP latent space. A traditional Transformer-based auto-encoder is employed to generate motion sequences. TEMOS [30] addresses stochastic motion generation using a VAE structure, aligning the embedding space of motion and text via KL-divergence between their latent distributions. Other works focus on employing various modern generation methods for improved conditional generation. For instance, T2M-GPT [31] utilizes Vector Quantised Variational AutoEncoder (VQ-VAE) to encode motion sequences into discrete tokens, enabling GPT-like autoregressive generation and training through next-token prediction, with text embedding serving as prior. MotionDiffuse [32] incorporates Denoising Diffusion Probabilistic Models (DDPM) into the task of text-motion generation, greatly enhancing the diversity and fidelity of generated sequences.
@@
In recent years, some works [88, 33, 34, 54, 36] aim for more general text-motion models capable of seamlessly handling various text-motion tasks simultaneously. For example, TM2T [33] establishes a bi-modal mutual mapping between texts and tokenized human motion using autoregressive neural machine translators (NMT), effectively handling both text-to-motion and motion-to-text tasks. MotionGPT [34] utilizes VQ-VAE to create a motion tokenizer and vocabulary, and then perform pretraining on both motion and text in a unified manner, by treating human motion as a foreign language. Additionally, some works explore various text-motion tasks beyond generation, including motion question answering [38], text-based motion editing [35, 36, 37], text-motion retrieval [89, 90], etc.
II-C State Space Model
@@
State space models (SSMs) are a series of sequential models renowned for their computation and memory efficiency and the ability to model long-term dependencies. The pioneering work, S4 [91], first proposed applying HiPPO [92] initialization to enable SSMs to maintain long-range memory. Subsequent studies [93, 94, 95, 96, 97, 98, 99] have followed this direction, further improving the space structure and network architecture of S4.
@@
Recently, Mamba [57, 100] introduces an input-dependent selection mechanism into SSMs, demonstrating linear time efficiency in long-sequence modeling and achieving outstanding performance across various sequential tasks. The model has been adapted to diverse tasks, including image restoration [101], image segmentation [102, 103, 104, 105, 106], point cloud [107], video understanding [108, 109, 110], pan-sharpening [111], graph analysis [112, 113], multimodal learning [58, 59]. There’re also some efforts [114, 115] aiming to establish a universal visual backbone based on Mamba by sequentializing images using patch-based methods akin to ViT. In a more recent study, Motion Mamba [116] presents a motion generation model that incorporates the Mamba block into a denoising U-Net architecture, effectively enhancing motion consistency across frames. In contrast to our method which deeply integrates text control into the selection mechanism, their model’s selective state space component remains unimodal, employed solely for spatial-temporal feature extraction of the motion sequence, while textual information is integrated during the diffusion process via a conditional denoiser.
@@
Figure 2: Dataset statistics of BABEL-Grounding. ‘Frame Number’ refers to the length of motion sequences. ‘Text Query Length’ denotes the length of textual annotations in the data. ‘Grounded Length Ratio’ indicates the ratio of the length of temporal segments corresponding to each text query to the total length of the sequence. ‘Segment Counts per Query’ refers to the number of temporal segments corresponding to each text query.
III BABEL-Grounding Dataset
The task of THMG involves determining the start and end timestamps of all the segments in the motion sequence that align with the given textual description. Our objective is to construct a dataset where each text query depicting specific human actions can be mapped to one or more temporal segments within the motion sequence. To this end, we construct the BABEL-Grounding dataset based on BABEL [60]. BABEL-Grounding provides detailed textual descriptions of human actions alongside their corresponding temporal segments in untrimmed motion sequences, with a total of 5,339 sequences with 21,307 text-segments annotations. Each sequence averages 743 frames, with the ground-truth temporal segments having an average frame count of 112, indicating their sparse distribution within the lengthy motion sequences. Figure 2 illustrates the dataset statistics of BABEL-Grounding. As depicted, the BABEL-Grounding dataset contains comprehensive textual descriptions for motion sequences of diverse lengths. Each text query may correspond to multiple temporal segments, and these segments are sparsely distributed throughout the entire motion sequence, inherently posing a challenge for motion grounding task.
Below, we provide a concise overview of the construction process for BABEL-Grounding annotations. The original BABEL dataset provides dense temporal annotations by labeling each temporal segment with corresponding actions. However, its textual annotations consist of simple categorical phrases rather than detailed and comprehensive sentences, and its data structure is not directly applicable to the THMG task. Therefore, a set of manually-crafted rules has been employed to augment the dataset. Figure 3 provides a schematic illustration of the data annotation pipeline and highlights the distinction between BABEL and BABEL-Grounding datasets. Details will be elaborated below.
III-A Textual augmentation
Utilizing external annotations
BABEL is built upon the AMASS motion capture database [1], while HumanML3D [24] offers detailed textual annotations at the sequence level for the AMASS database. For each entry in the BABEL dataset with overly simplistic annotations, its corresponding entry in HumanML3D is located by the sequence ID. When the sequence-level HumanML3D annotations contain the phrase from the BABEL annotations, human annotators will manually verify their correspondence and supplement the BABEL annotations with detailed textual descriptions from HumanML3D.
Template-based augmentation
As HumanML3D’s sequence-level annotations only partially cover the items in BABEL, many low-quality annotations remain unaddressed, particularly those with only one-word labels. To address this gap, we’ve manually crafted templates to enhance them. For example, ‘take’ is expanded to ‘take something with his hands’, ‘stir’ becomes ‘stir something with his hands in a circular motion’, and ‘place’ becomes ‘place the object at a specific location’. This substitution process enriches the overly simplistic annotations, rendering them more comprehensible for the model. The final textual annotations are further refined by GPT-3.5-turbo to make them more fluent, complete and diverse.
Figure 3: An illustration of the data augmentation pipeline, highlighting the differences between the original BABEL annotations and the BABEL-Grounding annotations.
III-B Temporal Augmentation
Time windows merging
one-to-many mapping
III-C Annotation Quality
III-C1 Automatic Evaluation
Figure 4: Evaluation results of annotation quality. $d=\frac{l}{3}$ measures the shifting offset relative to the time window length $l$. BABEL-Grounding achieves comparable matching accuracy to BABEL at zero offset, while demonstrating greater sensitivity to temporal shifts due to its richer information in text annotations.
To comprehensively evaluate the annotation quality of the proposed BABEL-Grounding dataset, we harness the power of Multimodal Large Language Model (MLLM) and design a protocol to assess both textual and temporal annotations. We select the open-source Qwen2.5-VL-72B-Instruct [117] for its superior video understanding capabilities, running at BF16 precision on two NVIDIA A100 GPUs.
For each data item, we first obtain the rendered 3D mesh video of the motion and segment the video according to the annotated time window to produce a clip. Next, we feed the clip along with its corresponding textual description into the MLLM, which outputs a ‘yes’ or ‘no’ judgment on whether the motion in the clip matches the text. We calculate the proportion of ‘yes’ responses as a matching accuracy indicator.
To evaluate the quality of our temporal annotations, we design a method based on time window shifting. Specifically, the annotated time windows are shifted by fractions of the window length $l$, i.e. ($\pm\frac{l}{3}$, $\pm\frac{2l}{3}$, $\pm l$), and then the matching accuracy with the text are computed using the same approach described above. The resulting accuracy curve, spanning from l-l to +l+l, is plotted in Figure 4, with the curve for the original BABEL also presented for comparison. As expected, maximum alignment is observed at zero offset, with accuracy steadily declining as the displacement increases. This indicates that shifting our annotated time window causes a mismatch between the text and the motion within the window, validating the precision of our temporal annotations.
Furthermore, compared to the fully manual annotations from the BABEL dataset, our automatic augmentation approach achieves annotation quality that is nearly equivalent to human-level standards (75.3% for BABEL and 74.9% for BABEL-Grounding at zero offset). We also note that as the shift distance increases, the accuracy gap between BABEL and BABEL-Grounding widens. This disparity arises because BABEL provides short, concise and sometimes ambiguous phrases, while our dataset uses detailed, complete sentences with richer information. Consequently, our text-motion alignment is more sensitive to temporal mismatches, as reflected in the steeper decline in matching accuracy when the time windows are shifted.
Figure 5: Results of dataset evaluation. The left side presents the frequency of scores ranging from 1 to 5, while the right side presents the cumulative percentages.
Score Statistics Mean Std Median Truncated mean
Text-motion Alignment 4.252 1.153 5.0 4.288
Segment Count Accuracy 4.413 1.115 5.0 4.514
Boundary Precision 4.223 1.063 5.0 4.270
TABLE I: Score statistics of dataset evaluation, including the mean, variance, median, and the truncated mean.
Figure 6: Interactive Gradio interface of the dataset evaluation system. Participants are presented with a video of a motion sequence together with its corresponding textual description and temporal segment annotations, and are required to provide rating scores for three questions.
III-C2 Manual Evaluation
we conducted a thorough manual evaluation to assess the quality of the dataset annotations. To be specific, we invited 25 participants and randomly sampled 1,000 instances from the dataset. Each participant was asked to evaluate 40 samples, rating the annotations with scores from 1 to 5 (where 1 is the worst and 5 is the best) across three aspects:
  • Text-Motion Alignment: semantic consistency between the motion and the textual description.
  • Segment Count Accuracy: correctness of the number of grounded temporal segments.
  • Boundary Precision: accuracy of the start and end time boundaries of each segment.
The evaluation interface was implemented by Gradio (see Fig. 6). Fig. 5 presents the distribution of ratings across the three aspects. As shown, approximately 80% of the samples received scores $\geq 4$, and about 90% of the samples received scores $\geq 3$. Tab. I summarizes statistical measures including mean, variance, and median scores for each aspect. To further account for the potential influence of extreme raters, we additionally report a truncated mean by excluding the top-3 and bottom-3 participants in terms of their overall average scores. In summary, both the automatic MLLM-based evaluation and the above manual evaluation consistently confirm the quality of the textual and temporal annotations in our BABEL-Grounding dataset.
Figure 7: Left: overall architecture of our proposed model. Right: Illustration of TM-Mamba block. ‘Bidirectional SSM’ refers to the text-controlled selection mechanism demonstrated in Algorithm 2 with bidirectional modeling. Text embedding q denotes the CLIP embedding of the input text query.